OpenStudy (anonymous):

what is the y-coordinate of the point of intersection of the graphs of: x+2y=-6 and 2x+y=24?

myininaya (myininaya):

x=-6-2y x=24/2-y/2=12-y/2 -6-2y=12-y/2 -2y=18-y/2 -2y+y/2=18 -4y/2+y/2=18 -3y/2=18 y=18(-2/3) y=6(-2) y=-12